OpenStudy (anonymous):

(5-2x/6) x (24/10-4x)

OpenStudy (anonymous):

is that "x" mean multiply?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

start by simplifying terms

OpenStudy (anonymous):

ok so we have ... (5-2x) /6 multiplied by (24/10-4x). if we look at the numerator 5-2x and denominator 10-4x we see that they are multiples of each other right? so you can write ... 10-4x = 2(5-2x)...so in this case we get ... (5-2x)/6 * 24/2*(5-2x) and then 5-2x cancel out so we are left with 6*(24/2) = 6*12 = 72

OpenStudy (anonymous):

my mistake . we are left with (1/6) * (24/2) = (1/6)*12 = 2 sorry its not 72, its 2. i forgot that 6 was in the denominator

OpenStudy (anonymous):

I think there's some confusion here about how you've written the problem. [ 5 - (2x/6) ][ (24/10) - 4x ] is not [ (5-2x)/6 ][ 24/(10-4x) ] Which one of these did you mean?